8:30: we start setting up the system for calibrations

9:30: ready to close the loop on ARGOS

we acquire a Board setup: 28 Nov 14_argos_calibunit with FLAO BIN4

PBL with BCU running at 1kHz continues to crash, we try to disable the ASM diagnostics. 700Hz still OK, but at 800Hz it cashes again.

We safely set the framerate to 500Hz.

command to restart the ARGOS arbitrator: on DX-LALAS terminal type:

argos_dx_lalas_monit restart DX.AARB

We close the loop on TT FLAO BIN4, ARGOS 500Hz, TS on 10 modes, no disturb.

TN Notes
20141128_192808 G=1,0.5,0.5
20141128_192951

We acquire the slopes_offset in closed loop with Stage X = -71.0 and

Piezo X Y
Blue -0.85 -0.86
Yellow -1.10 0.30
Red 0.90 1.40

We now move the Stage X to -71.1mm

TN Notes
20141128_203451 G=1,0.5,0.5
20141128_3547

We measure the new position for the piezos:

Piezo X Y
Blue -1.41 -0.25
Yellow -1.10 0.95
Red 0.33 -1.59

The jitter loop is stable, we load the disturb = 1.6", putting the gain to 0.5 or 0.6 makes the jitter to oscillate.

We create a fake REC and IM for the piezo to periscope offload: fake_imPP and we test the offload procedure.

New disturb considering the no RR case

Since the disturb files created the 13th of November were just copy-pasted from FLAO calibration directories they take into account the double reflection on the ASM. We produce new disturb taking into account this factor.

B Tag
1.2" 20141128_120000
1.6" 20141128_160000

We close the loop loading the 1.2" disturb.

TN Notes
20141128_224450 gain adjustment
20141128_224621 G=2,0.5,0.5
20141128_224705
20141128_224747
20141128_224953 OL
20141128_225057

-- MarcoBonaglia - 15 Dec 2014
Topic revision: r2 - 17 Dec 2014, MarcoBonaglia
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback